# High-Temperature Quantum Anomalous Hall Effect in Buckled Honeycomb Antiferromagnets

### v1

@@ Abstract
For the buckled honeycomb structure, see Fig. 1(a) for a side view, applying a perpendicular electric field induces a potential difference between the spin-up and the spin-down sublattices. This induces a shift of charges between the two kinds of sites. We approximate the multi-orbital Hubbard model by a generalized Kondo lattice model. This is well-justified if the charge fluctuations, i.e., the number of holes or double occupancies, do not exceed one per site. The generalized Kondo lattice model is given by (cf. Fig. 1)
@@
where $c^{{\dagger}}_{i\alpha}$ is the fermion creation operator at the site $i$ with the $z$-component of the spin $\alpha=\uparrow$ or $\downarrow$. The first term is the NN hopping and the second term is the Hund coupling between the electron spin $\vec{s}_{i}$ and the localized spin $\vec{S}_{i}$ with the spin quantum number $S=\mathcal{S}_{\rm tot}-1/2$. We treat one orbital as representative of the charge motion, while the others represent the localized spin. The third term in Eq. (2) is the Hubbard interaction with $n^{{\vphantom{\dagger}}}_{i\alpha}:=c^{{\dagger}}_{i\alpha}c^{{\vphantom{\dagger}}}_{i\alpha}$, and the fourth term is the Heisenberg interaction between the NN spins. We always count the Heisenberg interaction on each lattice bond only once. The fifth term is the staggered sublattice potential giving the onsite energies $\delta$ and $\delta$ to the two sublattices of the honeycomb structure. Figure 1(b) illustrates these different terms for the special case of $\mathcal{S}_{\rm tot}=3/2$.
@@
The number of methods which allow for a reliable investigation of the topological properties of the strongly correlated Hamiltonian (2) at finite temperature are rare. We employ the dynamical mean-field theory (DMFT) [28] as an established method for strongly correlated systems and use the exact diagonalization (ED) as the impurity solver [28, 29]. We specifically use the real-space realization of the DMFT [30] providing access to the bulk and the edge properties on equal footing [31]. Technical details can be found in Supplemental Materials [32]. We will compare the results for the different number of bath sites in the ED calculations.
@@
where $\boldsymbol{G}^{{\vphantom{\dagger}}}_{\alpha}\equiv\boldsymbol{G}^{{\vphantom{\dagger}}}_{\alpha}({\rm i}\omega_{n},\vec{k})$, $\epsilon_{\mu\nu\rho}$ is the totally antisymmetric tensor, and summations are implied over the indices $\mu$, $\nu$, and $\rho$, each of which runs over $\mathrm{i}\omega_{n}$, $k_{x}$, and $k_{y}$ [32].
@@
Figure 2: Phase diagram of temperature vs the alternating sublattice potential controlled by a perpendicular electric field. The colormap represents the value of the Hall conductance $\sigma_{yx}=-\sigma_{xy}$ given by Eq. (High-Temperature Quantum Anomalous Hall Effect in Buckled Honeycomb Antiferromagnets). The Néel temperature $T_{\rm N}$ and the crossover quantization temperature $T_{q}$ are specified. The Hall conductance takes the quantized value $e^{2}/h$ with an error less than $\%1$ below $T_{q}$, characterizing the antiferromagnetic Chern insulator (AFCI). The gray dotted lines separate the metallic region from the insulating regions. The metallic region shrinks rapidly to the two quantum critical points as $T\to 0$. The results are for the model parameters $S=1/2$, $U=12t$, $J_{\rm H}=0.2U$, $J=4t^{2}/\Delta_{0}=0.2\bar{7}t$, and $\lambda_{\rm SO}=0.2t$. We recall that the localized spin $S=1/2$ in Eq. (2) corresponds to the total spin $\mathcal{S}_{\rm tot}=S+1/2=1$ in Eq. (1). The number of bath sites $n_{b}=5$ is used in the ED impurity solver.
@@
For small values of $\delta$ the system is a paramagnetic Mott insulator at high and an AFI at low $T$, which is the well-known physics of the Heisenberg model [34]. Upon increasing $\delta$, the Néel temperature first increases. This is due to the increase in the effective Heisenberg interaction $J_{\rm eff}=4t^{2}\Delta_{0}/(\Delta_{0}^{2}-4\delta^{2})$ between the itinerant spins. For $\delta\gtrsim 5t$ the charge fluctuations start to contribute to the low-energy physics causing a drop in $T_{\rm N}$. For large values of $\delta$ the lower-energy sublattice is almost doubly occupied and the higher-energy sublattice is almost empty of electrons. The itinerant electrons show a very weak local magnetization due to the van Vleck mechanism [35, 32]. Thus, the itinerant electrons are in the band insulator phase above $T_{\rm N}$.
@@
The intermediate values of $\delta$ are of particular interest because the Hall conductance becomes non-zero and approaches the quantized value $e^{2}/h$ as the temperature is lowered. For $\delta=7t$ we plot the local magnetizations $m=|\langle s_{i}^{z}\rangle|$ and $M=|\langle S_{i}^{z}\rangle|$ in Fig. 3(a) and the Hall conductance vs $T$ in Fig. 3(b). Figure 3(c) represents the local spectral function averaged over the two sites in the unit cell for the spin component $\alpha$ near the Fermi energy $\omega=0$ at different temperatures. The data in Fig. 3 are for $n_{b}=6$ bath sites except for the filled gray squares at selective temperatures in panels (a) and (b) which are for $n_{b}=7$. The results for $n_{b}=5$, not included in the figure, also nicely match the results for $n_{b}=6$ and 7. This corroborates the accuracy of the results.
@@
Interestingly, despite the very different model parameters used in Fig. 6 one can hardly see any change in $T_{q}$. The quantization temperature is solely determined by the spin-orbit coupling and the hopping parameter. $T_{q}$ increases linearly with $\lambda_{\rm SO}$ and saturates for large values of $\lambda_{\rm SO}$. Although the spin-orbit coupling is highly compound-specific, one usually expect $\lambda_{\rm SO}<0.15t$. In this region, the data can nicely be described by the linear fit $T_{q}=0.92\lambda_{\rm SO}-0.003t$ shown in Fig. 6 as a black dashed line. This relation suggests a quantization temperature of the order of room temperature for $\lambda_{\rm SO}\sim 30$ meV and $t\sim 500$ meV, which are typical values for heavy transition-metal elements.
@@
Compounds Ba2NiTeO6 and Sr3CaOs2O9 are known buckled honeycomb antiferromagnets with the buckling heights of $5.3$ Å [39, 40] and $3.3$ Å [41], respectively. The latter compound shows a Néel AF order with the high Néel temperature $T_{\rm N}\sim 385$ K and is a promising candidate to host a high-$T$ AFCI. Growing $[111]$-oriented bilayers of perovskite heavy transition-metal oxides is another route to a buckled honeycomb antiferromagnet [42] and a high-$T$ AFCI state. Our findings open up a new path to the realization of the quantum anomalous Hall effect at high temperatures.

### v2

@@ Abstract
For the buckled honeycomb structure, see Fig. 1(a) for a side view, applying a perpendicular electric field induces a potential difference between the spin-up and the spin-down sublattices. This induces a shift of charges between the two kinds of sites. The description of the system requires the Heisenberg model (1) to be extended by the charge degrees of freedom, and the multi-orbital Hubbard model would be the natural choice. However, a reliable analysis of the topological properties of a strongly correlated multi-orbital model at finite temperature is a too ambitious numerical challenge. For this reason, we follow the idea proposed in Ref. [14] which was successful in the description of different compounds [12, 13]. We approximate the multi-orbital Hubbard model by a generalized Kondo lattice model. This is well-justified if the charge fluctuations, i.e., the number of holes or double occupancies, do not exceed one per site. The generalized Kondo lattice model is given by (cf. Fig. 1)
@@
where $c^{{\dagger}}_{i\alpha}$ is the fermion creation operator at the site $i$ with the $z$-component of the spin $\alpha=\uparrow$ or $\downarrow$. The first term is the NN hopping and the second term is the Hund coupling between the electron spin $\vec{s}_{i}$ and the localized spin $\vec{S}_{i}$ with the spin quantum number $S=\mathcal{S}_{\rm tot}-1/2$. We treat one orbital as representative of the charge motion, while the others represent the localized spin. We emphasize that the Kondo lattice approximation in Eq. (2) does not mean that we distinguish between the different orbitals of the same shape. The itinerant orbital stands as a representative for all the orbitals [14, 12, 13]. This restricts the possible amount of charge transfer between the higher and the lower energy sublattices. But, it is very well justified for the values of $\delta$ accessible via realistic strengths of electric fields, see below and also the ionicity discussed in Supplemental Materials [27].
The third term in Eq. (2) is the Hubbard interaction with $n^{{\vphantom{\dagger}}}_{i\alpha}:=c^{{\dagger}}_{i\alpha}c^{{\vphantom{\dagger}}}_{i\alpha}$, and the fourth term is the Heisenberg interaction between the NN spins. We always count the Heisenberg interaction on each lattice bond only once. The fifth term is the staggered sublattice potential giving the onsite energies $\delta$ and $\delta$ to the two sublattices of the honeycomb structure. Figure 1(b) illustrates these different terms for the special case of $\mathcal{S}_{\rm tot}=3/2$.
@@
The number of methods which allow for a reliable investigation of the topological properties of the strongly correlated Hamiltonian (2) at finite temperature are rare. We employ the dynamical mean-field theory (DMFT) [31] as an established method for strongly correlated systems and use the exact diagonalization (ED) as the impurity solver [31, 32]. We specifically use the real-space realization of the DMFT [33, 34, 35] providing access to the bulk and the edge properties on equal footing [36]. The lattice model is mapped to a set of effective Anderson-Kondo impurity models, which treat the localized spin quantum mechanically [12, 13]. The method takes the local charge and spin quantum fluctuations fully into account. The local quantum fluctuations leading to a frequency-dependent self-energy are essential for the description of the paramagnetic Mott insulator at finite temperature. The non-local Heisenberg interactions are replaced by an effective AF field via a mean-field decoupling [37]. The strength of the effective AF field is determined self-consistently during the DMFT loop [12, 13]. Further technical details can be found in Supplemental Materials [27]. Previous studies of similar interacting topological models by different methods indicate only minor effects to the phase boundaries due to the non-local quantum fluctuations, neglected in our investigation [38, 39, 40, 41, 42]. We will compare the results for the different number of bath sites in the ED calculations.
@@
[…]
@@
Figure 2: Phase diagram of temperature vs the alternating sublattice potential controlled by a perpendicular electric field. The colormap represents the value of the Hall conductance $\sigma_{yx}=-\sigma_{xy}$ given by Eq. (High-Temperature Quantum Anomalous Hall Effect in Buckled Honeycomb Antiferromagnets). The Néel temperature $T_{\rm N}$ and the crossover quantization temperature $T_{q}$ are specified. The Hall conductance takes the quantized value $e^{2}/h$ with an error less than $\%1$ below $T_{q}$, characterizing the antiferromagnetic Chern insulator (AFCI). The gray dotted lines separate the metallic region from the insulating regions. The metallic region shrinks rapidly to the two quantum critical points as $T\to 0$. The mean-field Néel temperature $\tilde{T}_{\rm N}$ of the effective low-energy spin model valid for small values of $\delta$, see the main text, is also included. The results are for the model parameters $S=1/2$, $U=12t$, $J_{\rm H}=0.2U$, $J=4t^{2}/\Delta_{0}=0.2\bar{7}t$, and $\lambda_{\rm SO}=0.2t$. We recall that the localized spin $S=1/2$ in Eq. (2) corresponds to the total spin $\mathcal{S}_{\rm tot}=S+1/2=1$ in Eq. (1). The number of bath sites $n_{b}=5$ is used in the ED impurity solver.
@@
For small values of $\delta$ the system is a paramagnetic Mott insulator at high and an AFI at low $T$, which is the well-known physics of the Heisenberg model [46]. Upon increasing $\delta$, the Néel temperature first increases. This is due to the increase in the effective Heisenberg interaction $J_{\rm eff}=4t^{2}\Delta_{0}/(\Delta_{0}^{2}-4\delta^{2})$ between the itinerant spins. The mean-field Néel temperature $\tilde{T}_{\rm N}$ of the effective low-energy spin model [27] nicely describes the Néel temperature $T_{\rm N}$ of the generalized Kondo lattice model up to $\delta\simeq 3t$. We attribute the small deviation to the contributions beyond the second order perturbation theory used in the derivation of the effective spin model. For $\delta\gtrsim 5t$ the charge fluctuations start to play a qualitative role in the low-energy physics causing a drop in $T_{\rm N}$. This behavior is reminiscent of the Néel temperature of the Hubbard model as the Hubbard interaction is reduced to lower values than in the Mott regime [47]. For large values of $\delta$ the lower-energy sublattice is almost doubly occupied and the higher-energy sublattice is almost empty of electrons. The itinerant electrons show a very weak local magnetization due to the van Vleck mechanism [48, 27]. Thus, the itinerant electrons are in the band insulator phase above $T_{\rm N}$.
The intermediate values of $\delta$ are of particular interest because the Hall conductance becomes non-zero and approaches the quantized value $e^{2}/h$ as the temperature is lowered. For $\delta=7t$ we plot the local magnetizations $m=|\langle s_{i}^{z}\rangle|$ and $M=|\langle S_{i}^{z}\rangle|$ in Fig. 3(a) and the Hall conductance vs $T$ in Fig. 3(b). Figure 3(c) represents the local spectral function averaged over the two sites in the unit cell for the spin component $\alpha$ near the Fermi energy $\omega=0$ at different temperatures. The finite spectral weight developing at the Fermi energy at the paramagnetic high temperatures signals a metallic state. An intermediate metallic phase separating the band and the Mott insulators is already reported for the ionic Hubbard model [49, 50]. Our results suggest that such a metallic phase appears also in multi-orbital systems at paramagnetic high temperatures. At low temperatures, the system is expected to be an insulator [51, 52, 53].
The data in Fig. 3 are for $n_{b}=6$ bath sites except for the filled gray squares at selective temperatures in panels (a) and (b) which are for $n_{b}=7$. The results for $n_{b}=5$, not included in the figure, also nicely match the results for $n_{b}=6$ and 7. This corroborates the accuracy of the results.
@@
Interestingly, despite the very different model parameters used in Fig. 6 one can hardly see any change in $T_{q}$. The quantization temperature is solely determined by the spin-orbit coupling and the hopping parameter. $T_{q}$ increases linearly with $\lambda_{\rm SO}$ and saturates for large values of $\lambda_{\rm SO}$. Although the spin-orbit coupling is highly compound-specific, one usually expect $\lambda_{\rm SO}<0.15t$. In this region, the data can nicely be described by the linear fit $T_{q}=0.92\lambda_{\rm SO}-0.003t$ shown in Fig. 6 as a black dashed line. This relation allows us to estimate the scale of the quantization temperature based on the knowledge of the spin-orbit coupling and the hopping parameter.
The precise determination of the strength of the spin-orbit coupling $\lambda_{\rm SO}$ in a material is challenging. However, it is usually estimated to be of the order of a few meV for $3d$, around $15$ meV for $4d$, and around $40$ meV for $5d$ transition-metal elements [54]. Note that the numbers refer to the Kane-Mele-type of the spin-orbit coupling. For comparison, the precise value of $\lambda_{\rm SO}$ for silicene, germanene, and stanene are computed to be $0.77$, $8.9$, and $12$ meV, respectively [55]. Assuming $t\sim 200$ meV for $3d$ transition-metal compounds one finds that it is possible to realize the AFCI in these materials although with the small quantization temperature $T_{q}\sim 10$ K. This estimate nicely agrees with first principles calculations predicting the AFCI in CrO with a small charge gap of about $1$ meV at zero temperature [22]. For the more extended $4d$ orbital with the typical hopping parameter $t\sim 400$ meV we find $T_{q}\sim 100$ K. Assuming $t\sim 700$ meV for $5d$ materials, our results unveil the possible realization of the AFCI at room temperature.
@@
Compounds Ba2NiTeO6 and Sr3CaOs2O9 are known buckled honeycomb antiferromagnets with the buckling heights of $5.3$ Å [59, 60] and $3.3$ Å [61], respectively. The latter compound shows a Néel AF order with the high Néel temperature $T_{\rm N}\sim 385$ K and is a promising candidate to host a high-$T$ AFCI. It should be noted that Sr3CaOs2O9 is not anisotropic enough to be considered effectively two-dimensional and thin-film growth would be needed. It is also conceivable to strain-engineer the films to enhance the in-plane and further suppress the out-of-plane couplings to put the system closer to the two-dimensional limit. Growing $[111]$-oriented bilayers of perovskite heavy transition-metal oxides is another route to a buckled honeycomb antiferromagnet [62] and a high-$T$ AFCI state. The possible realization of an AFCI in $[111]$-oriented perovskite bilayers has been previously predicted through studies of static exchange AF fields and first-principles calculations at zero temperature [63]. Our analysis going beyond static mean-field approaches and addressing the topological properties at finite temperatures explicitly unveils that AF Mott insulators can host a quantum anomalous Hall effect that can persist up to high temperatures.